UW CHEMISTRY NMR

Announcments & Equipment Logs

ALERT ! AV500 UPDATE ON PROBEHEAD – PLEASE READ

Tuning and Matching glitch on 1H channel :

The matching knob on 1H channel has lost its traction since it is mostly disengaged from the rod that rotates the matching Capacitor inside the probe.  We need to apply a minimally invasive remedy to fix this and we will do this on Monday, the 28th.  But the probehead has to come out of the magnet for the same to be completed.

The saving grace is that the match capacitor is at such a position that the matching efficiency is close to 90%.  

Recommendation:

Until we get this fixed next week, users are advised to only tune the probe (T – knob only) and proceed with their measurements. The tuning rod is working properly. 

NMR Facility Manager

1/25/19


Leave a Reply

Your email address will not be published. Required fields are marked *